# Stop Looking for “Important Tokens” in Multimodal Language Models: Duplication Matters More

### v1

@@ Abstract
Vision tokens in multimodal large language models often dominate huge computational overhead due to their excessive length compared to linguistic modality. Abundant recent methods aim to solve this problem with token pruning, which first defines an importance criterion for tokens and then prunes the unimportant vision tokens during inference. However, in this paper, we show that the importance is not an ideal indicator to decide whether a token should be pruned. Surprisingly, it usually results in inferior performance than random token pruning and leading to incompatibility to efficient attention computation operators. Instead, we propose DART (Duplication-Aware Reduction of Tokens), which prunes tokens based on its duplication with other tokens, leading to significant and training-free acceleration. Concretely, DART selects a small subset of pivot tokens and then retains the tokens with low duplication to the pivots, ensuring minimal information loss during token pruning. Experiments demonstrate that DART can prune 88.9% vision tokens while maintaining comparable performance, leading to a 1.99$\times$ and 2.99$\times$ speed-up in total time and prefilling stage, respectively, with good compatibility to efficient attention operators.
@@
Zichen Wen1,2, Yifeng Gao1, Shaobo Wang1, Junyuan Zhang2, Qintong Zhang2,4, Weijia Li3,2, Conghui He2222Corresponding authors., Linfeng Zhang1222Corresponding authors. 1Shanghai Jiao Tong University 2Shanghai AI Laboratory 3Sun Yat-sen University 4Peking University zichen.wen@outlook.com, heconghui@pjlab.org.cn, zhanglinfeng@sjtu.edu.cn
@@ 1 Introduction
Multimodal large language models (MLLMs) exhibit remarkable capabilities across a diverse range of multimodal tasks, including image captioning, visual question answering (VQA), video understanding (Wang et al., 2024b), and multimodal reasoning (Wang et al., 2024c).
@@
However, such impressive performance is always accompanied by huge computation costs, which are mainly caused by massive vision tokens in the input data, especially for high-resolution images (Li et al., 2024b) and multi-frame video (Tang et al., 2023), leading to challenges in their applications.
To solve this problem, abundant recent methods introduce token pruning to remove the vision tokens in a training-free manner, which usually first defines the importance score of each token, and then prunes the most unimportant tokens during the inference phrase (Chen et al., 2024; Zhang et al., 2024b; Liu et al., 2024e). The key to a token pruning method is the definition of the importance of vision tokens, where most existing methods are based on the attention scores between vision-only tokens and vision-language tokens. However, this paper argues that these importance-based methods have several serious problems.
@@
(III) Bias in token positions: As claimed by abundant recent works (Endo et al., 2024; Zhang et al., 2024a) and shown in Figure 1, attention scores have position bias, where the tokens are positionally close to the last token tend to have a higher attention score, making attention score does not truly reveal the value of this token.
(IV) Significant accuracy drop: Although the aforementioned three problems have reminded us the ineffectiveness of importance-based token pruning, however, it is still extremely surprising to find that some influential importance-based token pruning methods show inferior accuracy than random token pruning, (i.e., randomly selecting the tokens for pruning), as shown in Figure 2.
@@ 2 Related Work
2.1 Multimodal Large Language Models
Multimodal large language models (MLLMs) (Liu et al., 2024b; Li et al., 2023a; Zhu et al., 2023; Liu et al., 2024d) achieve remarkable performance in tasks involving images, videos, and multimodal reasoning by integrating vision and text processing. However, processing visual data poses computational challenges due to the redundancy and low information density of high-resolution tokens (Liang et al., 2022) and the quadratic scaling of attention mechanisms (Vaswani et al., 2017). For instance, models like LLaVA (Liu et al., 2023) and mini-Gemini-HD (Li et al., 2024b) encode high-resolution images into thousands of tokens, while video-based models such as VideoLLaVA (Lin et al., 2023) and VideoPoet (Kondratyuk et al., 2023) require even more tokens to process multiple frames. These challenges highlight the need for efficient token representations and longer context lengths to enable scalability. Recent advancements, such as Gemini (Team et al., 2023) and LWM (Liu et al., 2024a), address these issues by optimizing token efficiency and extending context length, paving the way for more scalable and effective MLLMs.
2.2 Visual Token Compression
Visual tokens often exceed text tokens by tens to hundreds of times, with visual signals being more spatially redundant compared to information dense text (Marr, 2010). Various methods have been proposed to address this issue. For instance, LLaMA-VID (Li et al., 2024a) uses a Q-Former with context tokens, and DeCo (Yao et al., 2024a) applies adaptive pooling to downsample visual tokens at the patch level. However, these approaches require modifying model components and additional training, increasing computational and training costs. ToMe (Bolya et al., 2023) reduces tokens without training by adding a token merge module to ViTs (Alexey, 2020), but this disrupts early cross-modal interactions in language models (Xing et al., 2024). FastV (Chen et al., 2024) selects important visual tokens using attention scores, while SparseVLM (Zhang et al., 2024b) incorporates text guidance via cross-modal attention. However, these methods forgo Flash-Attention (Dao et al., 2022; Dao, 2024) and primarily focus on token importance, overlooking the impact of token duplication. In our work, we maintain hardware acceleration compatibility, including Flash Attention, while focusing on another critical factor, token duplication, for efficient token reduction.
@@ 3.1 Preliminary
Architecture of MLLM. The architecture of Multimodal Large Language Models (MLLMs) typically comprises three core components: a visual encoder, a modality projector, and a language model (LLM). Given an image $I$, the visual encoder and a subsequent learnable MLP are used to encode $I$ into a set of visual tokens $e_{v}$. These visual tokens $e_{v}$ are then concatenated with text tokens $e_{t}$ encoded from text prompt $p_{t}$, forming the input for the LLM. The LLM decodes the output tokens $y$ sequentially, which can be formulated as: $y_{i}=f(I,p_{t},y_{0},y_{1},\cdots,y_{i-1}).$
@@ 3.4 Theoretical Analysis
SparseVLM (2024.10) 57.6 62.5 53.7 1721 83.6 69.1 75.6 56.1 50.5 292 96.3%
PDrop (2024.10) 57.1 63.2 56.8 1766 82.3 68.8 75.1 56.1 51.1 290 96.7%
@@
SparseVLM (2024.10) 56.0 60.0 51.1 1696 80.5 67.1 73.8 54.9 51.4 280 93.8%
PDrop (2024.10) 56.0 61.1 56.6 1644 82.3 68.3 72.9 55.1 51.0 287 95.1%
@@
SparseVLM (2024.10) 52.7 56.2 46.1 1505 75.1 62.2 68.2 51.8 50.1 180 84.6%
PDrop (2024.10) 41.9 33.3 50.5 1092 55.9 68.6 69.2 45.9 50.7 250 78.1%
@@
LLaVA-Next-7B Upper Bound, 2880 Tokens  (100%)
Vanilla 64.2 67.4 60.6 1851 86.5 70.1 81.8 64.9 57.6 517 100%
[…]
FastV (ECCV24) 55.9 61.6 51.9 1661 71.7 62.8 71.9 55.7 53.1 374 86.4%
HiRED (AAAI25) 59.3 64.2 55.9 1690 83.3 66.7 75.7 58.8 54.2 404 91.8%
SparseVLM (2024.10) 56.1 60.6 54.5 1533 82.4 66.1 71.5 58.4 52.0 270 85.9%
PDrop (2024.10) 56.4 63.4 56.2 1663 77.6 67.5 73.5 54.4 54.1 259 86.8%
MustDrop (2024.11) 57.3 62.8 55.1 1641 82.1 68.0 73.7 59.9 54.0 382 90.4%
FasterVLM (2024.12) 56.9 61.6 53.5 1701 83.6 66.5 74.0 56.5 52.6 401 89.8%
GlobalCom2 (2025.01) 57.1 61.8 53.4 1698 83.8 67.4 76.7 57.2 54.6 375 90.3%
DART (Ours) 61.7 65.3 58.2 1710 84.1 68.4 79.1 58.7 56.1 406 93.9%
Table 1: Comparative experiments on image understanding. In all experiments for DART, tokens are pruned after the second layer with $8$ pivot tokens. The pivot tokens are selected based on the maximum K-norm.
4 Experiments
Experiment Setting. We conduct experiments on over four MLLMs across ten image-based and four video-based benchmarks. For details on implementation, please refer to Appendix B.
@@ 4.1 Main Results
Image understanding task. The results presented in Table 1 highlight the exceptional performance of DART, across a diverse range of image understanding tasks under varying vision token configurations. We can observe that (i) with only $192$ tokens retained ($\downarrow 66.7\%$), DART achieves an impressive average performance of $98.8\%$, substantially outperforming the second-best method, MustDrop by $\mathbf{1.6\%}$. (ii) This trend becomes even more pronounced under more aggressive reduction ratios, with DART outperforming by $\mathbf{2.2\%}$ while retaining only $64$ tokens. (iii) Moreover, DART scales seamlessly to more advanced models, as evidenced by its performance on the LLaVA-Next-7B backbone, where it achieves an average score of $\mathbf{93.9\%}$ with only $11.1\%$ of visual tokens retained, outperforming all competing methods by a significant margin. These results highlight DART’s exceptional efficiency in leveraging limited visual tokens while preserving critical information, showcasing its robust performance across diverse tasks and seamless adaptability to various model architectures.
@@ 5.1 Efficiency Analysis
As shown in Table 2, we compare the total inference time, prefill time, FLOPs, and KV cache memory of multiple methods. (i) DART achieves a 2.99$\times$ speedup in the prefill phase and a 1.99$\times$ speedup in the entire inference, while its performance on POPE degrades by less than 3% compared to the vanilla model. (ii) Further analysis reveals that although the FLOPs reduction of different methods is similar, their actual inference speeds vary significantly. For instance, SparseVLM increases FLOPs by only 2.8% compared to DART, but its speedup drops by 21.6%, demonstrating that relying solely on FLOPs to measure acceleration effectiveness lacks practical significance. (iii) We further evaluate the performance-latency trade-off of each method based on actual latency. As illustrated in Figure 4, when considering the balance between latency and performance, some existing methods even underperform random visual token retention. We argue that SparseVLM and MustDrop suffer from significant speed degradation due to their sequential multi-stage token processing. For FastV, its reliance on biased attention scores for “important” token selection results in slightly inferior performance. In contrast, DART seamlessly integrates Flash Attention and incurs less than 0.08s overhead for token reduction, achieving a better trade-off between performance and speed.
@@ 5.3 Influence from Choice of the Pruned Layer and the Number of Pivot Tokens
We explore the impact of layer on model performance. As expected, pruning deeper layers yields performance closer to the vanilla model but increases latency, as shown in Figure 5. However, we observe two intriguing findings: (i) Pruning at layers 10, 15, and 20 surprisingly outperforms the vanilla model (Fig. 5(a)), consistent with Fig. 1, suggesting that removing duplicate tokens may reduce hallucinations in MLLMs on the POPE. (ii) At deeper layers (e.g., 15, 20), the latency-minimizing points correspond to pruning all vision tokens, yet performance drops only by $\mathbf{0.1\%}\mathbf{\sim}\mathbf{1.6\%}$. This highlights a modality imbalance in MLLMs, indicating underutilization of the visual modality.
Figure 6: Impact of the number of pivot tokens.
Furthermore, we delved into the impact of the number of pivot tokens on performance. As depicted in Figure 6, choosing either an insufficient or an excessive number of pivot tokens leads to suboptimal outcomes. When a limited number of pivot tokens (e.g., one or two), the lack of diversity among these tokens may impede their ability to comprehensively represent the entire feature space. In contrast, when an overly large number of pivot tokens, for example, 20 or more, are chosen, the majority of retained visual tokens tend to be pivot tokens. In extreme cases, our approach starts to resemble the importance-based method, where pivot tokens essentially transform into important tokens, overlooking the impact of duplication factors.
@@ 6 Conclusion
The pursuit of efficient token reduction in MLLMs has traditionally centered on token “importance” often measured by attention scores, sometimes suffering from worse performance than random token pruning. This study introduces DART, which focuses on the duplication of tokens, aiming to remove the tokens that are similar to others, leading to a superior balance between performance and latency in 13 benchmarks and 5 MLLMs (Tab. 1244, and Fig. 4). Our exploration yields surprising insights: multiple distinct sets of retained tokens, sharing less than 50% overlap, can deliver comparable and good performance (§5.2). Additionally, token pruning shows potential to mitigate hallucinations (§5.3). These observations highlight the limitations of importance-based methods and may offer valuable insights into the influence of vision tokens in MLLMs.
@@ 7 Limitations
One of the limitations of our work is that it cannot be applied to black-box models like the GPT (e.g. GPT 3.5 and more advanced versions) and Claude series, as we are unable to access their encoded tokens during the inference process. Additionally, our proposed DART exhibits slightly degraded performance when dealing with models that encode tokens with high information density. For instance, Qwen2-VL employs token merging during training, and MiniCPM-V-2.6 utilizes learnable queries in its Resampler module to map variable-length patch features into shorter feature representations. In such cases, removing a single visual token results in a greater loss of information, thereby impacting the effectiveness of our approach. Despite this, DART is still able to maintain 97.0% and 92.9% of its performance while reducing visual tokens by 66.7%, as demonstrated in Tables 4, 4.

### v2

@@ Abstract
Vision tokens in multimodal large language models often dominate huge computational overhead due to their excessive length compared to linguistic modality. Abundant recent methods aim to solve this problem with token pruning, which first defines an importance criterion for tokens and then prunes the unimportant vision tokens during inference. However, in this paper, we show that the importance is not an ideal indicator to decide whether a token should be pruned. Surprisingly, it usually results in inferior performance than random token pruning and leading to incompatibility to efficient attention computation operators. Instead, we propose DART (Duplication-Aware Reduction of Tokens), which prunes tokens based on its duplication with other tokens, leading to significant and training-free acceleration. Concretely, DART selects a small subset of pivot tokens and then retains the tokens with low duplication to the pivots, ensuring minimal information loss during token pruning. Experiments demonstrate that DART can prune 88.9% vision tokens while maintaining comparable performance, leading to a 1.99$\times$ and 2.99$\times$ speed-up in total time and prefilling stage, respectively, with good compatibility to efficient attention operators  [^1] Our code is available at https://github.com/ZichenWen1/DART.
@@
Zichen Wen1,2  Yifeng Gao1  Shaobo Wang1  Junyuan Zhang2  Qintong Zhang2,4 Weijia Li3,2Conghui He2222Corresponding authors.Linfeng Zhang1222Corresponding authors. 1Shanghai Jiao Tong University  2Shanghai AI Laboratory 3Sun Yat-sen University  4Peking University zichen.wen@outlook.com, heconghui@pjlab.org.cn, zhanglinfeng@sjtu.edu.cn
@@ 1 Introduction
Multimodal large language models (MLLMs) exhibit remarkable capabilities across a diverse range of multimodal tasks, including image captioning, visual question answering (VQA), video understanding (Wang et al., 2024b), and multimodal reasoning (Wang et al., 2024c; Kang et al., 2025).
@@
However, such impressive performance is always accompanied by huge computation costs, which are mainly caused by massive vision tokens in the input data, especially for high-resolution images (Li et al., 2024d) and multi-frame video (Tang et al., 2023), leading to challenges in their applications.
To solve this problem, abundant recent methods introduce token pruning to remove the vision tokens in a training-free manner, which usually first defines the importance score of each token, and then prunes the most unimportant tokens during the inference phrase (Chen et al., 2024; Zhang et al., 2024c; Liu et al., 2024e). The key to a token pruning method is the definition of the importance of vision tokens, where most existing methods are based on the attention scores between vision-only tokens and vision-language tokens. However, this paper argues that these importance-based methods have several serious problems.
@@
(III) Bias in token positions: As claimed by abundant recent works (Endo et al., 2024; Zhang et al., 2024b) and shown in Figure 1, attention scores have position bias, where the tokens are positionally close to the last token tend to have a higher attention score, making attention score does not truly reveal the value of this token.
(IV) Significant accuracy drop: Although the aforementioned three problems have reminded us of the ineffectiveness of importance-based token pruning, however, it is still extremely surprising to find that some influential importance-based token pruning methods show inferior accuracy than random token pruning, (i.e., randomly selecting the tokens for pruning), as shown in Figure 2.
@@ 2 Related Work
Multimodal Large Language Models Multimodal large language models (MLLMs) (Liu et al., 2024b; Li et al., 2023a; Zhu et al., 2023; Liu et al., 2024d) excel at image, video, and multimodal reasoning by integrating vision and text (Zhang et al., 2024a). However, visual data processing is costly due to redundancy, low information density (Liang et al., 2022; Liu et al., 2025b), and the quadratic cost of attention (Vaswani et al., 2017). For instance, models like LLaVA (Liu et al., 2023) and mini-Gemini-HD (Li et al., 2024d) encode high-resolution images into thousands of tokens, while video models like VideoLLaVA (Lin et al., 2023) and VideoPoet (Kondratyuk et al., 2023) handle even more tokens across frames. These challenges highlight the need for efficient token representations and longer context. Recent work like Gemini (Team et al., 2023) and LWM (Liu et al., 2024a) addresses this by improving token efficiency and extending context, enabling more scalable MLLMs.
Visual Token Compression Visual tokens often outnumber text tokens by tens to hundreds of times, as visual signals are more spatially redundant than information-dense text (Marr, 2010). LLaMA-VID (Li et al., 2024c) employs a Q-Former with context tokens, and DeCo (Yao et al., 2024a) uses adaptive pooling. DTMFormer (Wang et al., 2024d) improves ViTs’ efficiency in medical image segmentation by merging redundant tokens during training. MADTP (Cao et al., 2024) reduces computation by aligning cross-modal features and pruning tokens. However, these require modifying components and additional training. ToMe (Bolya et al., 2023) merges tokens without training but disrupts cross-modal interactions (Xing et al., 2024). FastV (Chen et al., 2024) selects via attention scores, while SparseVLM (Zhang et al., 2024c) uses text guidance. Yet, these forgo Flash-Attention (Dao et al., 2022; Dao, 2024), neglecting token duplication. We preserve hardware acceleration (i.e., Flash-Attention) and target duplication for efficient token reduction.
@@ 3.1 Preliminary
Architecture of MLLM. The architecture of Multimodal Large Language Models (MLLMs) typically comprises three core components: a visual encoder, a modality projector, and a language model (LLM). Given an image $I$, the visual encoder and a subsequent learnable MLP are used to encode $I$ into a set of visual tokens $e_{v}$. These visual tokens $e_{v}$ are then concatenated with text tokens $e_{t}$ encoded from the text prompt $p_{t}$, forming the input for the LLM. The LLM decodes the output tokens $y$ sequentially, which can be formulated as: $y_{i}=f(I,p_{t},y_{0},y_{1},\cdots,y_{i-1}).$
@@ 3.4 Theoretical Analysis
This provides a theoretical guarantee that DART preserves model output within a controllable bound, thereby supporting the trustworthiness and robustness of our method.
@@
SparseVLM (ICML25) 57.6 62.5 53.7 1721 83.6 69.1 75.6 56.1 50.5 292 96.3%
PDrop (CVPR25) 57.1 63.2 56.8 1766 82.3 68.8 75.1 56.1 51.1 290 96.7%
@@
SparseVLM (ICML25) 56.0 60.0 51.1 1696 80.5 67.1 73.8 54.9 51.4 280 93.8%
PDrop (CVPR25) 56.0 61.1 56.6 1644 82.3 68.3 72.9 55.1 51.0 287 95.1%
@@
SparseVLM (ICML25) 52.7 56.2 46.1 1505 75.1 62.2 68.2 51.8 50.1 180 84.6%
PDrop (CVPR25) 41.9 33.3 50.5 1092 55.9 68.6 69.2 45.9 50.7 250 78.1%
@@
DART (Ours) 57.1 64.7 56.7 1823 79.3 71.1 74.6 54.7 52.1 286 97.2%
Table 1: Comparative experiments on image understanding. In all experiments for DART, tokens are pruned after the second layer with $8$ pivot tokens. The pivot tokens are selected based on the maximum K-norm. DART indicates that DART is applied during the training stage of LLaVA-1.5-7B.
@@
4 Experiments
Experiment Setting. We conduct experiments on over four MLLMs across ten image-based and four video-based benchmarks. For details on implementation, please refer to Appendix C.
@@ 4.1 Main Results
[…]
LLaVA-Next-7B Upper Bound, 2880 Tokens  (100%)
Vanilla 64.2 67.4 60.6 1851 86.5 70.1 81.8 64.9 57.6 517 100.0%
[…]
FastV (ECCV24) 55.9 61.6 51.9 1661 71.7 62.8 71.9 55.7 53.1 374 86.4%
HiRED (AAAI25) 59.3 64.2 55.9 1690 83.3 66.7 75.7 58.8 54.2 404 91.8%
SparseVLM (ICML25) 56.1 60.6 54.5 1533 82.4 66.1 71.5 58.4 52.0 270 85.9%
PDrop (CVPR25) 56.4 63.4 56.2 1663 77.6 67.5 73.5 54.4 54.1 259 86.8%
MustDrop (2024.11) 57.3 62.8 55.1 1641 82.1 68.0 73.7 59.9 54.0 382 90.4%
FasterVLM (2024.12) 56.9 61.6 53.5 1701 83.6 66.5 74.0 56.5 52.6 401 89.8%
GlobalCom2 (2025.01) 57.1 61.8 53.4 1698 83.8 67.4 76.7 57.2 54.6 375 90.3%
DART (Ours) 61.7 65.3 58.2 1710 84.1 68.4 79.1 58.7 56.1 406 93.9%
Table 3: Comparative experiments are performed on LLaVA-Next-7B using the same settings as LLaVA-1.5-7B.
Image understanding task. The results presented in Tables 1 and 3 highlight DART’s exceptional performance across diverse image understanding tasks under varying token configurations. We observe that (i) with only $192$ tokens, DART achieves an impressive $98.8\%$ average performance, substantially outperforming second-best MustDrop by $\mathbf{1.6\%}$. (ii) This trend strengthens under aggressive reduction ratios, with DART leading by $\mathbf{2.2\%}$ using just $64$ tokens. (iii) Moreover, DART scales seamlessly to advanced and larger models like LLaVA-Next-7B and Qwen2-VL-72B (See Tab. 7), achieving $\mathbf{93.9\%}$ with only $11.1\%$ tokens, outperforming all competitors significantly. (iv) Inspired by (Wen et al., 2025), we apply DART during training. DART in Table 1 shows better performance-efficiency trade-offs, maintaining full performance with just 192 visual tokens, highlighting the strong adaptability of our method. These results demonstrate DART’s efficiency in leveraging limited tokens while preserving critical information, showcasing robust performance across tasks, model architectures, and model size. For more comparisons, please refer to Tables 55, and Appendix A.3.
@@ 5.1 Efficiency Analysis
As shown in Table 2, we compare the total inference time, prefill time, FLOPs, and KV cache memory of multiple methods. (i) DART achieves a 2.99$\times$ speedup in prefill and 1.99$\times$ speedup in inference, while its performance on POPE degrades by less than 3% versus the vanilla model. (ii) Analysis reveals although FLOPs reduction is similar across methods, their speeds vary significantly. For instance, SparseVLM increases FLOPs by 2.8% versus DART, but its speedup drops 21.6%, showing FLOPs alone poorly measure acceleration. (iii) We evaluate performance-latency trade-offs using actual latency. Figure 4 shows some methods underperform random token retention. SparseVLM and MustDrop suffer speed degradation from sequential token processing. FastV’s biased attention scores yield worse performance. In contrast, DART integrates Flash Attention with under 0.08s overhead, achieving better performance-speed balance.
@@ 5.3 Influence from Choice of the Pruned Layer and the Number of Pivot Tokens
Figure 5: Impact of the number of pivot tokens.
We explore the impact of layer on model performance. As expected, pruning deeper layers yields performance closer to the vanilla model but increases latency, as shown in Figure 6. However, we observe two intriguing findings: (i) Pruning at layers 10, 15, and 20 surprisingly outperforms the vanilla model (Fig. 6(a)), consistent with Fig. 1, suggesting that removing duplicate tokens may reduce hallucinations in MLLMs on the POPE. (ii) At deeper layers (e.g., 15, 20), the latency-minimizing points correspond to pruning all vision tokens, yet performance drops only by $\mathbf{0.1\%}\mathbf{\sim}\mathbf{1.6\%}$. This highlights a modality imbalance in MLLMs, indicating underutilization of the visual modality. Furthermore, we delved into the impact of the number of pivot tokens on performance. As depicted in Figure 5, choosing either an insufficient or an excessive number of pivot tokens leads to suboptimal outcomes. When a limited number of pivot tokens (e.g., one or two), the lack of diversity among these tokens may impede their ability to comprehensively represent the entire feature space. In contrast, when an overly large number of pivot tokens, for example, 20 or more, are chosen, the majority of retained visual tokens tend to be pivot tokens. In extreme cases, our approach starts to resemble the importance-based method, where pivot tokens essentially transform into important tokens, overlooking the impact of duplication factors.
@@ 6 Conclusion
The pursuit of efficient token reduction in MLLMs has traditionally focused on token “importance”, often measured by attention scores, but sometimes performs worse than random pruning. This study introduces DART, which targets token duplication, removing tokens similar to others and achieving better balance between performance and latency across multiple benchmarks and MLLMs (Tab. 1235579 and Fig. 4). Our exploration yields surprising insights: distinct retained token sets, with under 50% overlap, deliver similarly strong performance (§5.2). Moreover, token pruning may reduce hallucinations (§5.3). These findings expose limits of importance-based methods and offer insights into vision tokens in MLLMs.
@@ 7 Limitations
Similar to many other methods aimed at improving efficiency, such as network pruning, quantization, distillation, model merging, and speculative decoding, one of the limitations of our work is that it cannot be applied to black-box models like the GPT (e.g. GPT 3.5 and more advanced versions) and Claude series, as we are unable to access their encoded tokens during the inference process. Moreover, due to space limitations in the main text, we had to move some experimental results that we believe are particularly insightful and interesting to the appendix. These include, for example, our investigation of strategies for pivot token selection, a more detailed analysis of the impact of the number of pivot tokens, and validations of our method on larger-scale models, which may slightly affect the overall reading experience.
Appendix